# VeriLLM: A Lightweight Framework for Publicly Verifiable Decentralized Inference

### v2

@@ 1 Introduction
The recent surge of Large Language Models (LLMs) is driven by their broad utility in diverse domains, including software development [1] and creative content generation, resulting in widespread adoption across both industrial and academic settings [2, 3, 4]. However, this progress exposes a fundamental weakness in current AI infrastructure: the computational and model resources required to support large-scale inference are heavily centralized within a few dominant technology providers [5, 6, 7]. Such concentration leads to systemic risks, including privacy violations, restricted access, single points of failure, and monopolistic control over core AI capabilities [5, 8, 9, 10].
@@
To mitigate these risks, decentralized inference [11, 12, 13, 14, 15, 16, 17, 18] has emerged as a promising direction. By distributing computation across independent providers, decentralized systems aim to achieve openness, fault tolerance, and fairer access to AI resources [19, 20, 21, 22]. This paradigm enables verifiable collaboration among untrusted nodes and democratizes access to model execution, thereby improving transparency and resilience of large-scale AI deployments [23, 24, 20, 6].
@@
A central challenge in decentralized inference, however, is verifiable correctness. In a permissionless network, inference providers may be faulty, compromised, or economically motivated to skip computation. Without a mechanism to verify results, participants cannot ensure the integrity of model outputs [25, 26, 27, 28, 29]. This issue extends beyond adversarial behavior: providers often apply hidden optimizations—such as reduced numerical precision, pruned or distilled models, modified prompts, or compressed key–value caches—that degrade output fidelity [30, 31, 32, 33, 34, 35]. As a result, independent verification becomes an operational necessity for any decentralized inference infrastructure [36].
@@
Existing approaches to verification face critical efficiency and deployability limitations. Cryptographic proof systems, including zero-knowledge proofs (ZKPs), can formally attest to correctness but remain computationally prohibitive: current frameworks introduce overheads several orders of magnitude larger than native inference, particularly for transformer-scale models [37, 38, 39, 40]. Consensus-based schemes, in contrast, rely on redundant replication and majority voting or interactive verification games. These methods require strong synchrony or honest-majority assumptions and suffer from high latency and resource overhead [41, 42, 43]. Furthermore, both proof-based and replication-based approaches align poorly with modern blockchain environments, where on-chain verification of large ML proofs is expensive and deterministic numerical reproducibility across heterogeneous hardware remains unsolved.
@@
(1) Commitment of Intermediate States. During computation, all intermediate hidden states are committed to Merkle trees [44], with their root hashes recorded on-chain. These commitments yield tamper-evident logs, serving as binding evidence for subsequent verification and dispute resolution.
(2) Randomized Empirical Verification. A verifiable random function (VRF) [45] determines unpredictable verification indices. At these positions, any verifier can re-execute the corresponding computations off-chain and compare the locally obtained results with the committed Merkle values. This sampling-based mechanism substantially reduces verification overhead while maintaining probabilistic soundness.
(3) Dispute-Resolution Protocol. In the event of inconsistencies, any honest verifier may initiate a dispute-resolution procedure by submitting Merkle proofs that demonstrate mismatches between recomputed and committed states. When empirical verification is insufficient for resolution, the protocol escalates to zero-knowledge proofs [46, 47] to provide conclusive adjudication.
@@
Building on this foundation, we further optimize the verification path by exploiting the inherent computational structure of large language model (LLM) inference. A standard inference pipeline consists of a Prefill phase followed by an autoregressive Decoder phase, with the latter dominating total runtime [48]. We observe that when the complete output sequence is available to verifiers, the expensive Decoder phase need not be re-executed. Instead, verifiers can perform only the Prefill computation to validate correctness, as the subsequent token generation is a deterministic function of Prefill activations. This selective recomputation significantly reduces the verification overhead—approximately 1% of full inference cost—without compromising fidelity or verifiability.
Low verification cost alone is not sufficient in a permissionless network: validation must also be publicly auditable and resistant to non-participation. To achieve public verifiability and prevent lazy verification, VeriLLM augments the above optimization with a blockchain-assisted protocol that provides transparent coordination and accountable reporting, while preserving low on-chain cost. Intensive checks are executed off-chain by selected verifiers, and the blockchain performs lightweight consistency tests and resolves disputes. Sampling randomness is derived from on-chain verifiable random functions (VRFs) [45], which select unpredictable hidden-state positions for audit. For each sample, verifiers submit Merkle openings [44] with minimal recomputation to attest that the required steps were executed. A peer-consistency mechanism then compares submissions and rewards alignment with the distribution expected from honest computation, which discourages collusion and maintains liveness without heavy replication.
@@ 2 Design Goals and Threat Alignment
This section formalizes the security and system design goals of VeriLLM and establishes their alignment with the assumed adversarial capabilities. Our design aims to provide verifiable, tamper-evident, and publicly auditable large language model (LLM) inference in a decentralized environment without relying on trusted majorities or centralized coordination [26, 19, 27].
@@ 2.1 Adversarial Capabilities
We consider a permissionless network composed of three types of participants: Inferencers, Verifiers, and a Scheduler. Each participant may deviate from the prescribed protocol to gain computational or economic advantage [25, 41].
  • C1: Malicious Inferencer. An Inferencer may attempt to reduce computation cost by (i) running a smaller or quantized model, (ii) terminating early, or (iii) reusing cached or fabricated hidden states while pretending to produce valid outputs [38, 37].
  • C2: Lazy or Colluding Verifiers. A Verifier may skip recomputation and submit a verdict opportunistically, or collude with other verifiers or the Inferencer to falsely validate an incorrect computation [40, 29].
  • C3: Byzantine Scheduler. The Scheduler may selectively relay, alter, or censor messages; it may attempt to bias role assignment or falsify commitments to conceal misbehavior [30].
  • C4: Observation and Timing Adversary. The adversary can observe network-level messages and timing patterns in an attempt to infer whether a node is serving inference or verification, and may adapt behavior based on such information [49].
  • C5: On-chain Adversary. Attackers may attempt to bias or replay randomness, forge signatures, or submit inconsistent commitments to gain advantage in the commit–reveal process [45, 50].
@@ 2.2 Design Goals
G1. Public Verifiability. Any participant can verify inference correctness without relying on trust or replication. All hidden states and token outputs are bound by cryptographic commitments and, through a commit–then–sample procedure [39], any participant can open selected entries and check them against the model specification. This goal directly mitigates C1 and C2 by allowing verifiable detection of computation deviation or lazy verification. The verification process is auditable both off-chain (via hidden-state recomputation) and on-chain (via sampled scalar checks).
G2. Low Overhead. Verification overhead remains below 1% of the total inference cost. The scheme achieves this by limiting on-chain operations to sampled scalar openings rather than full tensor uploads, and by parallelizing verifier recomputation across layer segments [51]. This goal ensures practical deployability under realistic hardware and network conditions, addressing the performance constraints implied by C1 (economic motivation to cheat).
G3. One-Honest-Verifier Security. Correctness is guaranteed as long as at least one verifier executes the protocol honestly. Even if all but one verifier collude with the Inferencer or Scheduler, the presence of a single honest verifier ensures that inconsistencies or forged hidden states will be revealed either through the on-chain sampling phase or via zero-knowledge dispute proofs [40, 38]. This property directly counters C2 and C3 by removing the need for a majority or trusted quorum.
G4. Deterministic Accountability. Every message and hidden-state relay is traceable to a verifiable, signed commitment. Each segment boundary state is committed via a Merkle root [44] and signed by the responsible node. These signatures are registered on-chain, ensuring that any alteration, omission, or forgery by a malicious Scheduler (C3) or Inferencer (C1) is cryptographically detectable. This binding of state to identity guarantees tamper-evident trace integrity.
G5. Task-Type Indistinguishability. Workers cannot distinguish whether a job is an inference or a verification task. Inferencers and Verifiers receive identical message formats and API calls from the Scheduler, preventing selective deviation based on task type (mitigating C4). A node that attempts to skip verification will be caught with high probability during unpredictable on-chain sampling [11].
G6. Compatibility with Decentralized Deployment. The protocol operates robustly across heterogeneous GPUs and compute environments. VeriLLM tolerates bounded floating-point divergence by using calibrated statistical thresholds in hidden-state comparison [52]. This ensures that cross-device verification remains feasible and fair while preventing adversaries from exploiting hardware-level drift to mask malicious changes.
@@ 2.3 Threat–Goal Alignment
The combination of these goals forms a security envelope suited for large-scale decentralized inference. Public verifiability (G1) and deterministic accountability (G4) guarantee that every computation step is externally auditable. Low-overhead design (G2) ensures practical adoption without compromising latency or throughput. The one-honest-verifier guarantee (G3) provides strong correctness in adversarial environments, while task-type indistinguishability (G5) closes adaptive attack channels. Finally, cross-device compatibility (G6) grounds the protocol in realistic heterogeneous hardware deployments [51, 52]. Together, these principles ensure that VeriLLM achieves transparent, efficient, and verifiable LLM inference under adversarial conditions typical of open, decentralized systems.
@@ 3.1 Activation Sampling & Compact Commitment
Lightweight methods verify inference integrity by selectively committing to critical activation states. TOPLOC/INTELLECT [33] commits to the top-128 final-layer activations every 32 tokens, compressing data 1000× via polynomial congruence (258 bytes per commitment). However, its verification process requires a trusted third party to re-run the computation and provide a ground truth, which reintroduces a central trust assumption. TensorBlock [53] extends this to KV-cache sampling using deterministic token/layer selection, enabling validators to recompute partial trajectories from checkpoints. VeriSplit [54] supports private outsourcing through linear operator masking: clients add noise to inputs, workers return encrypted activations, and clients locally denoise results using precomputed terms, with Merkle trees enabling partial verification. However, their probabilistic security cannot guarantee the detection of sophisticated model tampering, and input privacy relies on noise injection techniques that are vulnerable to advanced reconstruction attacks.
@@ 3.2 Zero-Knowledge Machine Learning (ZKML)
Sumcheck-based Protocols: zkCNN [55] accelerates convolution verification by transforming operations to the frequency domain. zkLLM [39] reduces communication by 90% for non-arithmetic operations (e.g., Softmax) via the tlookup protocol and exploits translation invariance. QAP-based Protocols: vCNN [56] represents convolutions as Quadratic Polynomial Programs (QPP), reducing multiplicative gates exponentially. pvCNN [57] further optimizes tensor operations via Quadratic Matrix Programs (QMP) and batch verification. ZKML [38] offers modular circuit design with 43 pre-built layer circuits and automated optimization. zkGPT [40] leverages multi-thread parallelization for LLM inference-proof acceleration. Lookup arguments [58] enable matrix-level privacy with 30% faster proofs. VOLE-based Protocols: Mystique [59] achieves sublinear communication for matrix multiplication using sVOLE-based tensor compression. Recent frameworks [60] optimize nonlinear functions via table lookups and digital decomposition, reducing constraints by over 99%.
System & Hardware Optimization: zkLoRA [61] supports verifiable fine-tuning via incremental proofs for LoRA modules. EZKL [37] automates ZKP circuit generation from ONNX models. DeepProve [62] accelerates GPU-based ZKML (54–158× faster proofs via parallelized SNARKs). Sertn [63] employs optimistic ”Staked Deferred Proofs” with economic slashing for invalid results. However, these methods impose prohibitive prover costs (hours per inference), require trusted setups, and suffer from rigidity in supporting novel architectures due to circuit specialization.
@@ 3.3 Trusted Execution Environments (TEEs)
Atoma Network [64] combines TEE attestation with sampling-based consensus for deterministic execution. evML [65] uses TEE-based remote attestation (e.g., Google/Apple frameworks) with probabilistic auditing. nesa.ai [66] orchestrates heterogeneous TEEs for secure multi-party inference. Phala Network [67] extends this paradigm by providing a ”Confidential AI” cloud platform that enables standard Dockerized AI applications to run within GPU-equipped TEEs. However, they inherit vulnerabilities to side-channel exploits (e.g., cache-timing attacks), centralize trust in proprietary hardware (Intel SGX/ARM TrustZone), and exhibit limited auditability across heterogeneous environments.

### v3

@@ 1 Introduction
The recent surge of Large Language Models (LLMs) is driven by their broad utility in diverse domains, including software development [wang2025enhancingcodellmsreinforcement] and creative content generation, resulting in widespread adoption across both industrial and academic settings [brown2020language, openai2023gpt4, touvron2023llama]. However, this progress exposes a fundamental weakness in current AI infrastructure: the computational and model resources required to support large-scale inference are heavily centralized within a few dominant technology providers [bender2021dangers, zaharia2023accelerating, sevilla2022compute]. Such concentration leads to systemic risks, including privacy violations, restricted access, single points of failure, and monopolistic control over core AI capabilities [bender2021dangers, whittlestone2021sociotechnical, weidinger2023sociotechnical, hazan2023foundation].
@@
To mitigate these risks, decentralized inference [yuan2022decentralized, morpheus2024, gonka2024, awsneuron2024, mei2024fedmoe, menon2024ravnest, yin2024knowledge, qu2025prompt] has emerged as a promising direction. By distributing computation across independent providers, decentralized systems aim to achieve openness, fault tolerance, and fairer access to AI resources [allen2022decentralized, kuppuswamy2022decentralizedai, raj2023federated, he2021trustworthy]. This paradigm enables verifiable collaboration among untrusted nodes and democratizes access to model execution, thereby improving transparency and resilience of large-scale AI deployments [villalobos2022future, bai2023open, kuppuswamy2022decentralizedai, zaharia2023accelerating].
@@
A central challenge in decentralized inference, however, is verifiable correctness. In a permissionless network, inference providers may be faulty, compromised, or economically motivated to skip computation. Without a mechanism to verify results, participants cannot ensure the integrity of model outputs [castro1999practical, cachin2011introduction, zhang2020delphi, xu2021verifynet, zamani2023decentralized]. This issue extends beyond adversarial behavior: providers often apply hidden optimizations—such as reduced numerical precision, pruned or distilled models, modified prompts, or compressed key–value caches—that degrade output fidelity [ho2018security, wang2022publiccheck, scaramuzza2025zkm, ong2025toploc, wang2024modelcompression, xu2023compressprompt]. As a result, independent verification becomes an operational necessity for any decentralized inference infrastructure [huang2024trustworthiness].
@@
Existing approaches to verification face critical efficiency and deployability limitations. Cryptographic proof systems, including zero-knowledge proofs (ZKPs), can formally attest to correctness but remain computationally prohibitive: current frameworks introduce overheads several orders of magnitude larger than native inference, particularly for transformer-scale models [ezkl-github, chen2024zkml, sun2024zkllm, qu2025zkgpt]. Consensus-based schemes, in contrast, rely on redundant replication and majority voting or interactive verification games. These methods require strong synchrony or honest-majority assumptions and suffer from high latency and resource overhead [arun2025verde, mira-network, ambient-ai]. Furthermore, both proof-based and replication-based approaches align poorly with modern blockchain environments, where on-chain verification of large ML proofs is expensive and deterministic numerical reproducibility across heterogeneous hardware remains unsolved.
@@
(1) Commitment of Intermediate States. During computation, all intermediate hidden states are committed to Merkle trees [merkle1989certified], with their root hashes recorded on-chain. These commitments yield tamper-evident logs, serving as binding evidence for subsequent verification and dispute resolution.
(2) Randomized Empirical Verification. A verifiable random function (VRF) [micali1999verifiable] determines unpredictable verification indices. At these positions, any verifier can re-execute the corresponding computations off-chain and compare the locally obtained results with the committed Merkle values. This sampling-based mechanism substantially reduces verification overhead while maintaining probabilistic soundness.
(3) Dispute-Resolution Protocol. In the event of inconsistencies, any honest verifier may initiate a dispute-resolution procedure by submitting Merkle proofs that demonstrate mismatches between recomputed and committed states. When empirical verification is insufficient for resolution, the protocol escalates to zero-knowledge proofs [ben1988knowledge, groth2016size] to provide conclusive adjudication.
@@
Building on this foundation, we further optimize the verification path by exploiting the inherent computational structure of large language model (LLM) inference. A standard inference pipeline consists of a Prefill phase followed by an autoregressive Decoder phase, with the latter dominating total runtime [agrawal2023]. We observe that when the complete output sequence is available to verifiers, the expensive Decoder phase need not be re-executed. Instead, verifiers can perform only the Prefill computation to validate correctness, as the subsequent token generation is a deterministic function of Prefill activations. This selective recomputation significantly reduces the verification overhead—approximately 1% of full inference cost—without compromising fidelity or verifiability.
Low verification cost alone is not sufficient in a permissionless network: validation must also be publicly auditable and resistant to non-participation. To achieve public verifiability and prevent lazy verification, VeriLLM augments the above optimization with a blockchain-assisted protocol that provides transparent coordination and accountable reporting, while preserving low on-chain cost. Intensive checks are executed off-chain by selected verifiers, and the blockchain performs lightweight consistency tests and resolves disputes. Sampling randomness is derived from on-chain verifiable random functions (VRFs) [micali1999verifiable], which select unpredictable hidden-state positions for audit. For each sample, verifiers submit Merkle openings [merkle1989certified] with minimal recomputation to attest that the required steps were executed. A peer-consistency mechanism then compares submissions and rewards alignment with the distribution expected from honest computation, which discourages collusion and maintains liveness without heavy replication.
@@ 2 Design Goals and Threat Alignment
This section formalizes the security and system design goals of VeriLLM and establishes their alignment with the assumed adversarial capabilities. Our design aims to provide verifiable, tamper-evident, and publicly auditable large language model (LLM) inference in a decentralized environment without relying on trusted majorities or centralized coordination [cachin2011introduction, allen2022decentralized, zhang2020delphi].
@@ 2.1 Adversarial Capabilities
We consider a permissionless network composed of three types of participants: Inferencers, Verifiers, and a Scheduler. Each participant may deviate from the prescribed protocol to gain computational or economic advantage [castro1999practical, arun2025verde].
  • C1: Malicious Inferencer. An Inferencer may attempt to reduce computation cost by (i) running a smaller or quantized model, (ii) terminating early, or (iii) reusing cached or fabricated hidden states while pretending to produce valid outputs [chen2024zkml, ezkl-github].
  • C2: Lazy or Colluding Verifiers. A Verifier may skip recomputation and submit a verdict opportunistically, or collude with other verifiers or the Inferencer to falsely validate an incorrect computation [qu2025zkgpt, zamani2023decentralized].
  • C3: Byzantine Scheduler. The Scheduler may selectively relay, alter, or censor messages; it may attempt to bias role assignment or falsify commitments to conceal misbehavior [ho2018security].
  • C4: Observation and Timing Adversary. The adversary can observe network-level messages and timing patterns in an attempt to infer whether a node is serving inference or verification, and may adapt behavior based on such information [scion2015secure].
  • C5: On-chain Adversary. Attackers may attempt to bias or replay randomness, forge signatures, or submit inconsistent commitments to gain advantage in the commit–reveal process [micali1999verifiable, bonneau2015sok].
@@ 2.2 Design Goals
G1. Public Verifiability. Any participant can verify inference correctness without relying on trust or replication. All hidden states and token outputs are bound by cryptographic commitments and, through a commit–then–sample procedure [sun2024zkllm], any participant can open selected entries and check them against the model specification. This goal directly mitigates C1 and C2 by allowing verifiable detection of computation deviation or lazy verification. The verification process is auditable both off-chain (via hidden-state recomputation) and on-chain (via sampled scalar checks).
G2. Low Overhead. Verification overhead remains below 1% of the total inference cost. The scheme achieves this by limiting on-chain operations to sampled scalar openings rather than full tensor uploads, and by parallelizing verifier recomputation across layer segments [rajbhandari2022deep]. This goal ensures practical deployability under realistic hardware and network conditions, addressing the performance constraints implied by C1 (economic motivation to cheat).
G3. One-Honest-Verifier Security. Correctness is guaranteed as long as at least one verifier executes the protocol honestly. Even if all but one verifier collude with the Inferencer or Scheduler, the presence of a single honest verifier ensures that inconsistencies or forged hidden states will be revealed either through the on-chain sampling phase or via zero-knowledge dispute proofs [qu2025zkgpt, chen2024zkml]. This property directly counters C2 and C3 by removing the need for a majority or trusted quorum.
G4. Deterministic Accountability. Every message and hidden-state relay is traceable to a verifiable, signed commitment. Each segment boundary state is committed via a Merkle root [merkle1989certified] and signed by the responsible node. These signatures are registered on-chain, ensuring that any alteration, omission, or forgery by a malicious Scheduler (C3) or Inferencer (C1) is cryptographically detectable. This binding of state to identity guarantees tamper-evident trace integrity.
G5. Task-Type Indistinguishability. Workers cannot distinguish whether a job is an inference or a verification task. Inferencers and Verifiers receive identical message formats and API calls from the Scheduler, preventing selective deviation based on task type (mitigating C4). A node that attempts to skip verification will be caught with high probability during unpredictable on-chain sampling [yuan2022decentralized].
G6. Compatibility with Decentralized Deployment. The protocol operates robustly across heterogeneous GPUs and compute environments. VeriLLM tolerates bounded floating-point divergence by using calibrated statistical thresholds in hidden-state comparison [shoeybi2019megatron]. This ensures that cross-device verification remains feasible and fair while preventing adversaries from exploiting hardware-level drift to mask malicious changes.
@@ 2.3 Threat–Goal Alignment
The combination of these goals forms a security envelope suited for large-scale decentralized inference. Public verifiability (G1) and deterministic accountability (G4) guarantee that every computation step is externally auditable. Low-overhead design (G2) ensures practical adoption without compromising latency or throughput. The one-honest-verifier guarantee (G3) provides strong correctness in adversarial environments, while task-type indistinguishability (G5) closes adaptive attack channels. Finally, cross-device compatibility (G6) grounds the protocol in realistic heterogeneous hardware deployments [rajbhandari2022deep, shoeybi2019megatron]. Together, these principles ensure that VeriLLM achieves transparent, efficient, and verifiable LLM inference under adversarial conditions typical of open, decentralized systems.
@@ 3.1 Activation Sampling & Compact Commitment
Lightweight methods verify inference integrity by selectively committing to critical activation states. TOPLOC/INTELLECT [ong2025toploc] commits to the top-128 final-layer activations every 32 tokens, compressing data 1000× via polynomial congruence (258 bytes per commitment). However, its verification process requires a trusted third party to re-run the computation and provide a ground truth, which reintroduces a central trust assumption. TensorBlock [proof-of-cache-github] extends this to KV-cache sampling using deterministic token/layer selection, enabling validators to recompute partial trajectories from checkpoints. VeriSplit [zhang2024verisplit] supports private outsourcing through linear operator masking: clients add noise to inputs, workers return encrypted activations, and clients locally denoise results using precomputed terms, with Merkle trees enabling partial verification. However, their probabilistic security cannot guarantee the detection of sophisticated model tampering, and input privacy relies on noise injection techniques that are vulnerable to advanced reconstruction attacks.
@@ 3.2 Zero-Knowledge Machine Learning (ZKML)
Sumcheck-based Protocols: zkCNN [liu2021zkcnn] accelerates convolution verification by transforming operations to the frequency domain. zkLLM [sun2024zkllm] reduces communication by 90% for non-arithmetic operations (e.g., Softmax) via the tlookup protocol and exploits translation invariance. QAP-based Protocols: vCNN [lee2024vcnn] represents convolutions as Quadratic Polynomial Programs (QPP), reducing multiplicative gates exponentially. pvCNN [weng2023pvcnn] further optimizes tensor operations via Quadratic Matrix Programs (QMP) and batch verification. ZKML [chen2024zkml] offers modular circuit design with 43 pre-built layer circuits and automated optimization. zkGPT [qu2025zkgpt] leverages multi-thread parallelization for LLM inference-proof acceleration. Lookup arguments [campanelli2024lookup] enable matrix-level privacy with 30% faster proofs. VOLE-based Protocols: Mystique [weng2021mystique] achieves sublinear communication for matrix multiplication using sVOLE-based tensor compression. Recent frameworks [hao2024scalable] optimize nonlinear functions via table lookups and digital decomposition, reducing constraints by over 99%.
System & Hardware Optimization: zkLoRA [roy2025zklora] supports verifiable fine-tuning via incremental proofs for LoRA modules. EZKL [ezkl-github] automates ZKP circuit generation from ONNX models. DeepProve [deep-prove-github] accelerates GPU-based ZKML (54–158× faster proofs via parallelized SNARKs). Sertn [sertn-avs-github] employs optimistic ”Staked Deferred Proofs” with economic slashing for invalid results. However, these methods impose prohibitive prover costs (hours per inference), require trusted setups, and suffer from rigidity in supporting novel architectures due to circuit specialization.
@@ 3.3 Trusted Execution Environments (TEEs)
Atoma Network [atoma-infer-github] combines TEE attestation with sampling-based consensus for deterministic execution. evML [evml-github] uses TEE-based remote attestation (e.g., Google/Apple frameworks) with probabilistic auditing. nesa.ai [nesa-github] orchestrates heterogeneous TEEs for secure multi-party inference. Phala Network [phala2024] extends this paradigm by providing a ”Confidential AI” cloud platform that enables standard Dockerized AI applications to run within GPU-equipped TEEs. However, they inherit vulnerabilities to side-channel exploits (e.g., cache-timing attacks), centralize trust in proprietary hardware (Intel SGX/ARM TrustZone), and exhibit limited auditability across heterogeneous environments.